# Variational inequalities and smooth-fit principle for singular stochastic control problems in Hilbert spaces

### v1

@@ 3.1. Preliminary properties of $V$
    There exists $\hat{c}_{o}>0$ such that, for $p\geq 2$ and $\kappa_{2}>0$ as in Assumption 2.8,
    $-\kappa_{2}\leq V(x)\leq\hat{c}_{o}(1+|x|^{p}_{H});$
@@
    $V$ is semiconcave with semiconcavity constant $\hat{c}_{1}$; that is, there exists $\hat{c}_{1}>0$ such that
    $\lambda V(x)+(1-\lambda)V(y)-V(\lambda x+(1-\lambda)y)\leq\frac{\hat{c}_{1}}{2}\lambda(1-\lambda)|x-y|_{H}^{2},\ \ \forall x,y\in H,\ \lambda\in[0,1];$
@@
$V(x)$ $\leq\mathcal{J}(x;0)=\mathsf{E}\left[\int_{0}^{\infty}e^{-\rho t}{G}(X_{t}^{x,0})\mathrm{d}t\right]\leq c_{0}\int_{0}^{\infty}e^{-\rho t}\left(1+\mathsf{E}[|X_{t}^{x,0}|_{H}^{p}]\mathrm{d}t\right)$
$\leq{c_{0}}\left(\frac{1}{\rho}+2^{p-1}\int_{0}^{\infty}e^{-\rho t}\left(|e^{t\mathcal{A}}x|_{H}^{p}+\mathsf{E}[|W_{t}^{\sigma,\mathcal{A}}|_{H}^{p}]\right)\mathrm{d}t\right)$
$\leq{c_{0}}\left(\frac{1}{\rho}+2^{p-1}\int_{0}^{\infty}e^{-(\rho+p\delta)t}|x|_{H}^{p}\mathrm{d}t+2^{p-1}\frac{1}{\rho}\overline{c}_{p}\right)$
$={c_{0}}\left(\frac{1}{\rho}\big{(}1+2^{p-1}\overline{c}_{p}\big{)}+\frac{2^{p-1}}{\rho+p\delta}|x|_{H}^{p}\right),$
@@
Proof of (iii). It follows from the previous items and, e.g., [25, Cor. 2.4, p. 12].
@@
Proof of (v). This follows by Lemma (i)(ii). ∎
@@
(3.1) ^$\hat{\mathcal{I}}_{0}:=\left\{I=(\vartheta,\nu)\in\mathcal{I}_{0}:\ \mathsf{E}\bigg{[}\int_{0^{-}}^{\infty}e^{-\rho t}\mathrm{d}\nu_{t}\bigg{]}<\infty\right\}.$
@@ 3.2. Dynamic programming equation and viscosity solutions
where $\Theta$ is as in (2.15) and where we have defined the second-order differential operator $\mathcal{G}$, formally associated to the process $X^{0}$ and acting on sufficiently smooth functions $v:H\to\mathbb{R}$, as
(3.3) $[\mathcal{G}v](x):=\langle\mathcal{A}x,Dv(x)\rangle_{H}+\frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}v(x)\right].$
@@
Hence, in order to provide the definition of viscosity solution to (3.2), we first introduce the suitable class of test functions, and we then have a Dynkin’s formula for those functions (see Proposition 3.3 below). For a given but arbitrary $\widehat{C}>0$, let
$\mathcal{X}:=$$\Big{\{}\varphi\in C^{2}(H):\ D\varphi\in C(H;\mathcal{D}(\mathcal{A})),\ \ \sigma\sigma^{*}D^{2}\varphi\in C(H;\mathcal{L}_{1}(H)),\,\,|\varphi(x)|\leq\widehat{C}(1+|x|_{H}^{2}),$
(3.5) $\ \ \ \ \ \ \ \ \ |D\varphi(x)|_{H}+|\mathcal{A}D\varphi(x)|_{H}+\ |\sigma\sigma^{*}D^{2}\varphi(x)|_{\mathcal{L}_{1}(H)}\leq\widehat{C}\,\,\,\forall x\in H\Big{\}}.$
Remark 3.2.
In the context of classical stochastic control, the set $\mathcal{X}$, combined with the set of radial functions, is typically large enough to prove comparison results for viscosity solutions of the PDE (see [27, Ch. 3]), and therefore to characterize the value function of the control problem as the unique solution to the associated HJB equation.
In this paper, we do not address the (relevant) topic of uniqueness (see also Remark 3.9 below). However, we will use test functionw from the set $\mathcal{X}$ – of quadratic form – in Section 4 in order to prove our regularity results (see, in particular, the proof of Proposition 4.15).
In the rest of this paper, for any measurable function $f:H\to\mathbb{R}$, we set, for any $\tau\in\mathcal{T}$ and any $I\in\mathcal{I}$,
@@
We then have the next Dynkin’s formula for functions in $\mathcal{X}$. In the subsequent analysis we shall use that any $I\in\mathcal{I}$ is identified with an element $(\vartheta,\nu)\in\mathcal{I}_{0}$ (see Lemma 2.11 and the discussion afterwards) and that the optimization in (2.20) can be actually performed over $\hat{\mathcal{I}_{0}}$ (cf. (3.1)).
@@
Let $x\in H$, $\varphi\in\mathcal{X}$, $I\in\hat{\mathcal{I}_{0}}$, and $\tau\in\mathcal{T}$. Assume that
(3.7) $\mathsf{E}\left[\int_{0^{-}}^{\tau}e^{-\rho t}\left|\langle D\varphi(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\right|\mathrm{d}\nu_{t}\right]<\infty.$
Then, the following Dynkin’s formula holds true:
@@
Let $(\mathcal{A}_{n})_{n\in\mathbb{N}}$ be the Yosida approximants of $\mathcal{A}$ (see, e.g., [26, Eq. (3.7), Ch. II]), and let $X^{n;x,I}$ denote the solution to
@@
$[\mathcal{G}_{n}\varphi](x):=\langle\mathcal{A}_{n}x,D\varphi(x)\rangle+\frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}\varphi(x)\right].$
Thanks to [26, Lemma 3.4(ii), Ch. 2] and due to the fact that $\mathcal{A}$ is dissipative (cf. Assumption 2.1), we then have, as $n\uparrow\infty$,
(3.9) $\mathcal{A}_{n}y\to\mathcal{A}y\ \ \ \mbox{and}\ \ \ \sup_{n\in\mathbb{N}}|\mathcal{A}_{n}y|_{H}\leq|\mathcal{A}y|_{H}\quad\forall y\in\mathcal{D}(\mathcal{A}),$
@@
Moreover, for $p\geq 2$ as in Assumption 2.8 (see also (2.12)), by Lemma B.1(ii) (see also [27, Prop. 1.132] for the regular control case and references therein) it holds
(3.11) $\lim_{n\to\infty}\mathsf{E}[\sup_{t\in[0,T]}|X_{t}^{n;x,I}-X_{t}^{x,I}|^{p}]=0.$
The latter implies – up to passing to a subsequence – that $X_{t}^{n;x,I}\to X_{t}^{x,I}$ and with probability one as $n\to\infty$. Also, by [52, Thm. 27.2], for any $\varphi\in\mathcal{X}$, $\tau\in\mathcal{T}$, and for $T>0$, the following Dynkin’s formula holds true
$\mathsf{E}[e^{-\rho(\tau\wedge T)}\varphi(X^{n;x,I}_{\tau\wedge T})]=\varphi(x)+\mathsf{E}\bigg{[}\int_{0}^{\tau\wedge T}e^{-\rho t}[(\mathcal{G}_{n}-\rho)\varphi](X^{n;x,I}_{t})\mathrm{d}t\bigg{]}$
(3.12) $+\mathsf{E}\bigg{[}\int_{0^{-}}^{\tau\wedge T}e^{-\rho t}\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg{]}.$
As a matter of fact, the growth condition $|\sigma\sigma^{*}D^{2}\varphi(x)|_{\mathcal{L}_{1}(H)}\leq\widehat{C}$, for some $\widehat{C}>0$, as required for any test function $\varphi\in\mathcal{X}$ ensures that the local martingale term vanishes in expectation.
We now aim at taking $n\to\infty$ in (3.2). In this regard, the only convergences that require some attention are
(3.13) $\lim_{n\to\infty}\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}\langle X^{n;x,I}_{t},\mathcal{A}_{n}D\varphi(X^{n;x,I}_{t})\rangle_{H}\mathrm{d}t\right]=\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}\langle X^{x,I}_{t},{\mathcal{A}}D\varphi(X^{x,I}_{t})\rangle_{H}\mathrm{d}t\right]$
@@
(3.14) $\lim_{n\to\infty}\mathsf{E}\bigg{[}\int_{0^{-}}^{\tau\wedge T}e^{-\rho t}\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg{]}=\mathsf{E}\bigg{[}\int_{0^{-}}^{\tau\wedge T}e^{-\rho t}\langle D\varphi(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg{]}.$
All the other convergences can be performed by using arguments as in the proof of [27, Prop. 1.164] (in the case of regular controls). Assuming then that limits as $n\uparrow\infty$ can be indeed interchanged with expectations and integrals in (3.2) (we will verify later that (3.13) and (3.14) do hold), we find that
$\mathsf{E}\left[e^{-\rho(\tau\wedge T)}\varphi(X^{x,I}_{\tau\wedge T})\right]=\varphi(x)+\mathsf{E}\bigg{[}\int_{0}^{\tau\wedge T}e^{-\rho t}[(\mathcal{G}-\rho)\varphi](X^{x,I}_{t})\mathrm{d}t\bigg{]}$
$+\mathsf{E}\bigg{[}\int_{0^{-}}^{\tau\wedge T}e^{-\rho t}\langle D\varphi(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg{]},$
so that finally letting $T\uparrow\infty$ we obtain (3.8).
It thus remains to check the validity of (3.13) and (3.14). As for (3.14), recalling (3.2) one has
$\Big{|}\mathds{1}_{[0,\tau\wedge T]}(t)e^{-\rho t}\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\Big{|}\leq\widehat{C}|{\vartheta}_{t}|_{H}$
and, because $(\vartheta,\nu)\in\mathcal{I}_{0}$ (cf. (2.17)),
$\mathsf{E}\bigg{[}\int_{0^{-}}^{T}\Big{|}\mathds{1}_{[0,\tau\wedge T]}(t)e^{-\rho t}\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}|\,\mathrm{d}\nu_{t}\bigg{]}\leq\widehat{C}\mathsf{E}\bigg{[}\int_{0^{-}}^{T}|\vartheta_{t}|_{H}\,\mathrm{d}\nu_{t}\bigg{]}<\infty.$
Hence, the dominated convergence theorem implies (3.14), upon recalling that (up to a subsequence) $X_{t}^{n;x,I}\to X_{t}^{x,I}$ with probability one as $n\to\infty$.
@@
[…]
@@
Therefore, by using Hölder’s inequality, for a constant $C>0$ that may vary from line to line, we find
$\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}\Big{|}\langle X_{t}^{n;x,I},\mathcal{A}_{n}D\varphi(X_{t}^{n;x,I})\rangle_{H}-\langle X_{t}^{x,I},{\mathcal{A}}D\varphi(X^{x,I}_{t})\rangle_{H}\Big{|}\,\mathrm{d}t\right]$
$\leq C\left(\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}|X^{n;x,I}_{t}-X_{t}^{x,I}|_{H}^{2}\mathrm{d}t\right]\,\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}|D\varphi(X_{t}^{n;x,I})|^{2}_{\mathcal{D}({\mathcal{A}})}\mathrm{d}t\right]\right)^{1/2}$
$+C\left(\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}|X_{t}^{x,I}|_{H}^{2}\mathrm{d}t\right]\,\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}|D\varphi(X_{t}^{n;x,I})-D\varphi(X_{t}^{x,I})|^{2}_{\mathcal{D}({\mathcal{A}})}\mathrm{d}t\right]\right)^{1/2}$
$+\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}|X^{x,I}_{t}|_{H}\,\,|({\mathcal{A}}_{n}-{\mathcal{A}})D\varphi(X^{x,I}_{t})|_{H}\mathrm{d}t\right].$
We now verify that the dominated convergence theorem can be applied when taking limits as $n\uparrow\infty$ in order to show that the right-hand side of the last inequality converges to zero and thus (3.13) holds. We provide details only for the term
$\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}|D\varphi(X_{t}^{n;x,I})-D\varphi(X_{t}^{x,I})|^{2}_{\mathcal{D}({\mathcal{A}})}\mathrm{d}t\right]$
since all the others can be treated by similar arguments (thanks also to (3.11)). Recall (3.4) and (3.2). Then
$\Big{|}\mathds{1}(t)_{[0,\tau\wedge T]}e^{-\rho t}|D\varphi(X_{t}^{n;x,I})-D\varphi(X_{t}^{x,I})|^{2}_{\mathcal{D}({\mathcal{A}})}\Big{|}\leq\widehat{C}^{2},$
so that the dominated convergence theorem ensures that
$\lim_{n\uparrow\infty}\mathsf{E}\left[\int_{0}^{\tau\wedge T}e^{-\rho t}|D\varphi(X_{t}^{n;x,I})-D\varphi(X_{t}^{x,I})|^{2}_{\mathcal{D}({\mathcal{A}})}\mathrm{d}t\right]=0,$
upon recalling again that (up to a subsequence) $X_{t}^{n;x,I}\to X_{t}^{x,I}$ with probability one as $n\to\infty$. ∎
The next proposition provides an inequality for the cube of the norm on $H$. This will be exploited in the proof of the viscosity property of $V$ (see Theorem 3.7 below).

### v2

@@ 3.1. Preliminary properties of $V$
    There exists $\hat{c}_{o}>0$ such that, for $\kappa_{2}>0$ as in Assumption 2.7, -κ_2 V(x)≤^c_o(1+—x—^2_H);
@@
    $V$ is semiconcave with semiconcavity constant $\hat{c}_{1}$; that is, there exists $\hat{c}_{1}>0$ such that λV(x)+(1-λ) V(y) -V(λx+(1-λ) y)≤^c12λ(1-λ)—x-y—_H^2, x,yH, λ[0,1];
@@
$V(x)$ $\leq\mathcal{J}(x;0)=\mathsf{E}\left[\int_{0}^{\infty}e^{-\rho t}{G}(X_{t}^{x,0})\mathrm{d}t\right]\leq c_{0}\int_{0}^{\infty}e^{-\rho t}\left(1+\mathsf{E}[|X_{t}^{x,0}|_{H}^{2}]\mathrm{d}t\right)$
$\leq{c_{0}}\left(\frac{1}{\rho}+2\int_{0}^{\infty}e^{-\rho t}\left(|e^{t\mathcal{A}}x|_{H}^{2}+\mathsf{E}[|W_{t}^{\sigma,\mathcal{A}}|_{H}^{2}]\right)\mathrm{d}t\right)$
$\leq{c_{0}}\left(\frac{1}{\rho}+2\int_{0}^{\infty}e^{-(\rho+2\delta)t}|x|_{H}^{2}\mathrm{d}t+2\frac{1}{\rho}\overline{c}_{2}\right)$
$={c_{0}}\left(\frac{1}{\rho}\big(1+2\overline{c}_{2}\big)+\frac{2}{\rho+2\delta}|x|_{H}^{2}\right),$
@@
Proof of (iii). It follows from the previous items and, e.g., [26, Cor. 2.4, p. 12].
@@
Proof of (v). This follows by [48], end of page 265. ∎
@@
(3.1) $\hat{\mathcal{I}}_{0}:=\left\{I=(\vartheta,\nu)\in\mathcal{I}_{0}:\ \mathsf{E}\bigg[\int_{0^{-}}^{\infty}e^{-\rho t}\mathrm{d}\nu_{t}\bigg]<\infty\right\};$
that is, we have
(3.2) $V(x)=\hat{V}(x):=\inf_{(\vartheta,\nu)\in\hat{\mathcal{I}}_{0}}\mathcal{J}(x;I),\quad x\in H.$
@@ 3.2. Dynamic programming equation and viscosity solutions
where $\Theta$ is as in (2.17) and where we have defined the second-order differential operator $\mathcal{G}$, formally associated to the process $X^{x,0}$ and acting on sufficiently smooth functions $v:H\to\mathbb{R}$, as
(3.4) $\mathcal{G}v](x):=\langle\mathcal{A}x,Dv(x)\rangle_{H}+\frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}v(x)\right].$
@@
3.2.1. Dynkin’s formulae
In order to provide the definition of viscosity solution to (3.3), we first introduce the suitable class of test functions, and we then prove Dynkin’s formulae for those functions. Let us start by introducing the set of functions
$\mathcal{X}:=$ $\Big\{\varphi\in C^{2}(H):\ D\varphi\in C(H;\mathcal{D}(\mathcal{A})),\ \ \sigma\sigma^{*}D^{2}\varphi\in C(H;\mathcal{L}_{1}(H)),\$
(3.6) $\ \ \ \ \ \ \ \ \ \varphi(x)\leq\hat{C}(1+|x|_{H}^{2}),\ |D\varphi(x)|_{\mathcal{D}(A)}\leq\hat{C}(1+|x|_{H}),\ |\sigma\sigma^{*}D^{2}\varphi(x)|_{\mathcal{L}_{1}(H)}\leq\hat{C}\Big\}$.
In the subsequent analysis we shall use that any $I\in\mathcal{I}$ is identified with an element $(\vartheta,\nu)\in\mathcal{I}_{0}$ (see Lemma 2.10 and the discussion afterwards) and that the optimization in (2.22) can be actually performed over $\hat{\mathcal{I}}_{0}$ (cf. (3.1)). Also, in the rest of this paper, for any measurable function $f:H\to\mathbb{R}$, we set, for any $\tau\in\mathcal{T}$ and any $I\in\mathcal{I}$,
@@ 3.2.1. Dynkin’s formulae
We then have the next Dynkin formula for functions belonging to $\mathcal{X}$.
@@
Let $x\in H$, $\varphi\in\mathcal{X}$, $I\in\hat{\mathcal{I}}_{0}$. Let $\tau\in\mathcal{T}$ be bounded and such that there exists $M>0$ such that $\sup_{0\leq t\leq\tau}|X^{x,I}_{t}|\leq M$. Then, the following Dynkin formula holds true:
@@
Let $(\mathcal{A}_{n})_{n\in\mathbb{N}}$ be the Yosida approximants of $\mathcal{A}$ (see, e.g., [27, Eq. (3.7), Ch. II]), and let $X^{n;x,I}$ denote the solution to
@@
$\mathcal{G}_{n}\varphi](x):=\langle\mathcal{A}_{n}x,D\varphi(x)\rangle+\frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}\varphi(x)\right].$
Thanks to [27, Lemma 3.4(ii), Ch. 2] and due to the fact that $\mathcal{A}$ is dissipative (cf. Assumption 2.1), we then have,
(3.9) $\mathcal{A}_{n}y\to\mathcal{A}y\ \ \mbox{as}\ n\uparrow\infty\ \ \mbox{and}\ \ \ \sup_{n\in\mathbb{N}}|\mathcal{A}_{n}y|_{H}\leq|\mathcal{A}y|_{H}\quad\forall y\in\mathcal{D}(\mathcal{A}).$
@@
Recalling that $\tau$ is bounded, so that $\tau\leq T<\infty$ for some $T>0$, we may now use [54, Thm. 27.2], upon noticing that the stochastic integral
$\int_{0}^{\tau}\langle\mathcal{A}_{n}X_{t}^{n;x,I},D\varphi(X_{t}^{n;x,I})\rangle_{H}\mathrm{d}W_{t}$
vanishes in expectations by Lemma B.1(i) and by the growth assumption on $D\varphi$. We therefore obtain the following Dynkin’s formula :
(3.11) $\mathsf{E}[e^{-\rho\tau}\varphi(X^{n;x,I}_{\tau})]=\varphi(x)+\mathsf{E}\bigg[\int_{0}^{\tau}e^{-\rho t}\left[[(\mathcal{G}_{n}-\rho)\varphi](X^{n;x,I}_{t})\mathrm{d}t+\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\right]\bigg].$
We now aim at taking $n\to\infty$ in (3.11) to get the claim. By Lemma B.1(ii) (see also [28, Prop. 1.132] for the regular control case and references therein), it holds that $X_{t}^{n;x,I}\to X_{t}^{x,I}$ in probability for every $t\in[0,T]$ as $n\to\infty$. The only convergences that require some attention are
(3.12) $\lim_{n\to\infty}\mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}\langle X^{n;x,I}_{t},\mathcal{A}_{n}D\varphi(X^{n;x,I}_{t})\rangle_{H}\mathrm{d}t\right]=\mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}\langle X^{x,I}_{t},{\mathcal{A}}D\varphi(X^{x,I}_{t})\rangle_{H}\mathrm{d}t\right]$
@@
(3.13) $\lim_{n\to\infty}\mathsf{E}\bigg[\int_{0^{-}}^{\tau}e^{-\rho t}\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg]=\mathsf{E}\bigg[\int_{0^{-}}^{\tau}e^{-\rho t}\langle D\varphi(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg].$
All the other convergences can be performed by using arguments as in the proof of [28, Prop. 1.164] (in the case of regular controls).
As for (3.13), recalling (3.2.1) one has
$\Big|\mathds{1}_{[0,\tau]}(t)e^{-\rho t}\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\Big|\leq\widehat{C}(1+|X^{n;x,I}_{t}|_{H})|{\vartheta}_{t}|_{H}$
and because $\tau\in\mathcal{T}$ is bounded (say, by some $T>0$) we have for $q\in(1,2)$, upon using Hölder’s inequality,
$\mathsf{E}\bigg[\int_{0^{-}}^{T}\Big|\mathds{1}_{[0,\tau]}(t)e^{-\rho t}\langle D\varphi(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\Big|^{q}\,\mathrm{d}\nu_{t}\bigg]\leq\widehat{C}\mathsf{E}\bigg[\int_{0^{-}}^{T}(1+|X^{n;x,I}_{t}|^{q})|\vartheta_{t}|^{q}_{H}\,\mathrm{d}\nu_{t}\bigg],$
$\leq\widehat{C}\mathsf{E}\bigg[\sup_{0\leq t\leq T}(1+|X^{n;x,I}_{t}|^{q})\int_{0^{-}}^{T}|\vartheta_{t}|^{q}_{H}\mathrm{d}\nu_{t}\bigg]$
$\leq\widehat{C}\mathsf{E}\bigg[\sup_{0\leq t\leq T}(1+|X^{n;x,I}_{t}|^{\frac{2q}{2-q}})\bigg]^{\frac{2-q}{2}}\mathsf{E}\bigg[\int_{0^{-}}^{T}|\vartheta_{t}|^{2}_{H}\mathrm{d}\nu_{t}\bigg]^{\frac{q}{2}}<\infty$
where the finiteness of the last terms above is due to the fact that $(\vartheta,\nu)\in\mathcal{I}_{0}$ (cf. (2.19)) and Lemma B.1(i). Hence, Vitali’s convergence theorem implies (3.13), upon recalling that $X_{t}^{n;x,I}\to X_{t}^{x,I}$ in probability for every $t\in[0,T]$ as $n\to\infty$.
@@
$\leq|x_{n}-x|_{H}|{\mathcal{A}}_{n}y_{n}|_{H}+|x|_{H}|\mathcal{A}_{n}(y_{n}-y)|_{H}+|x|_{H}|({\mathcal{A}}_{n}-{\mathcal{A}})y|_{H}\vskip 6.0pt plus 2.0pt minus 2.0pt$
@@
Therefore, by using Hölder’s inequality and also that $\tau\in\mathcal{T}$ is bounded (say, by some $T>0$) and such that there exists $M>0$ for which $\sup_{0\leq\tau}|X^{x,I}_{t}|\leq M$, we find,
$\mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}\Big|\langle X_{t}^{n;x,I},\mathcal{A}_{n}D\varphi(X_{t}^{n;x,I})\rangle_{H}-\langle X_{t}^{x,I},{\mathcal{A}}D\varphi(X^{x,I}_{t})\rangle_{H}\Big|\,\mathrm{d}t\right]$
(3.15) $\leq\left(\mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}|X^{n;x,I}_{t}-X_{t}^{x,I}|_{H}^{2}\mathrm{d}t\right]\,\mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}|D\varphi(X_{t}^{n;x,I})|^{2}_{\mathcal{D}({\mathcal{A}})}\mathrm{d}t\right]\right)^{1/2}$
$+M\mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}|D\varphi(X_{t}^{n;x,I})-D\varphi(X_{t}^{x,I})|_{\mathcal{D}({\mathcal{A}})}\mathrm{d}t\right]$
$+M\mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}|({\mathcal{A}}_{n}-{\mathcal{A}})D\varphi(X^{x,I}_{t})|_{H}\mathrm{d}t\right].$
We now verify that limits as $n\uparrow\infty$ can be interchanged with the expectations in order to show that the right-hand side of (3.2.1) converges to zero and thus (3.12) holds.
By Lemma B.1(i) and the fact that that $X_{t}^{n;x,I}\to X_{t}^{x,I}$ in probability for every $t\in[0,T]$ when $n\to\infty$, it is not hard to be convinced that the first addend in the right-hand side of (3.2.1) converges to zero by the dominated convergence theorem.
Also, recalling (3.2.1) one has for a constant $C>0$
$\mathds{1}_{[0,\tau]}(t)e^{-\rho t}\Big|D\varphi(X^{n;x,I}_{t})-D\varphi(X^{x,I}_{t})\Big|^{2}_{\mathcal{D}(\mathcal{A})}\leq{C}(1+|X^{n;x,I}_{t}|^{2}_{H}+|X^{x,I}_{t}|^{2}_{H}).$
By Lemma B.1(i) the right-hand side of the previous equation are $\mathsf{P}\otimes\mathrm{d}t$-integrable over $\Omega\times[0,T]$, so that Vitali’s convergence theorem can be invoked in order to show that the second addend in the right-hand side of (3.2.1) convergence to zero, upon recalling that $D\varphi\in C(H;\mathcal{D}(\mathcal{A})$ and that, again, $X_{t}^{n;x,I}\to X_{t}^{x,I}$ in probability for every $t\in[0,T]$.
Similarly, also the last addend in the right-hand side of (3.2.1) can be shown to converge to zero by Vitali’s convergence theorem, upon noticing that by the right hand-side of (3.9) and (3.2.1) (see also (3.5)) one has, for some $C>0$,
$|({\mathcal{A}}_{n}-{\mathcal{A}})D\varphi(X^{x,I}_{t})|^{2}_{H}\leq 4|{\mathcal{A}}D\varphi(X^{x,I}_{t})|^{2}_{H}\leq{C}(1+|X^{x,I}_{t}|^{2}_{H}),$
with the last term above being $\mathsf{P}\otimes\mathrm{d}t$-integrable over $\Omega\times[0,T]$ due to Lemma B.1(i), and upon using also that $({\mathcal{A}}_{n}-{\mathcal{A}})y\to 0$ as $n\uparrow\infty$, for any $y\in\mathcal{D}(\mathcal{A})$ (cf. (3.9)). ∎
The next proposition provides an inequality of Dynkin type for radial functions. Let us introduce the set of smooth radial functions (centered at $z\in H$) as
(3.16) $\mathcal{R}^{z}:=\big\{g:H\to\mathbb{R}:\ g(y):=g_{0}(|y-z|_{H})\ \mbox{with}\ g_{0}\in C^{2}(\mathbb{R}^{+};\mathbb{R}),\ g^{\prime}_{0}\geq 0,\ g^{\prime}_{0}(0)=0\big\}.$
Notice that, if $g\in\mathcal{R}^{z}$, then $g\in C^{2}(H;\mathbb{R})$. However, since
(3.17) $Dg(y)=\begin{cases}\displaystyle{\frac{g_{0}^{\prime}(|y-z|_{H})}{{|y-z|_{H}}}(y-z),\quad\ \ \ \ \mbox{if}\ y\neq z,}\\ 0,\quad\quad\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \mbox{if}\ y=z,\end{cases}$
we do not have in general $Dg(y)\in\mathcal{D}(\mathcal{A})$, which prevents to use the Dynkin formula provided by Proposition 3.3. Nevertheless, the inequality provided in the new result will allow to include also functions of this type in the definition of viscosity solutions (see Definition 3.6).